OpenStudy (anonymous):

A sedan will drive ten miles north, 15 miles east, 13 miles south, 15 miles west, and 22 miles north. The sedan gets 20 miles per gallon, and there are five gallons of gas left in the tank. Compute the relative distance and the total distance. Then decide if there is enough fuel for the trip.

OpenStudy (anonymous):

@DemiUnicorn

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

Okay, you do the same thing as last time. You add 10 + 15 + 13 + 15 + 22 = 75. Since the Sedan get 20 miles per gallon and there is 5 gallons left in the tank, you multiply 5 and 20. So, 5 x 20 = 100. So, 100 - 75= 25. After making the trip you can still have enough gas to go another 25 miles. Meaning, you do have enough gas to make the trip.

OpenStudy (anonymous):

Write two different scenarios that are similar to the three previous problems above. Compute the relative distance and the total distance, and then decide if there is enough fuel for the trip. Review the rubric to understand how the problems will be evaluated.

OpenStudy (anonymous):

I think you can do that one on your own. It's really simple. Just follow the questions. To help, I'll give you the first one. A mustang will drive 5 miles south, 10 miles east, and 15 miles north. The mustang gets 15 miles per gallon and there are 3 gallons of gas in the car. So you add 5 + 10 + 15 = 30. Then 15 x 3 = 45. $5 - 30 = 15. So, after the trip you will still be able to travel another 15 miles, or there is one gallon of gas left in the tank. Now, you can do the other scenario.

OpenStudy (anonymous):

thanks I get it now id give you another medal if I could lol

OpenStudy (anonymous):

Lmao, you are very welcome.

OpenStudy (anonymous):

You can become a fan x0x

OpenStudy (anonymous):

will do

OpenStudy (anonymous):

I am